OpenStudy (kendricklamar2014):

Solve the Equation using addition or subtraction. Check your answer. Q - 11 = -9

OpenStudy (kendricklamar2014):

@Destinymasha @HelpBlahBlahBlah @jim_thompson5910

OpenStudy (helpblahblahblah):

okay so if im not mistaking this is how you would do the equation: Q- 11= -9 +11 +11 ------- ------ 0 2 so Q=2

OpenStudy (helpblahblahblah):

then to check your answer you would do 2-11= -9

OpenStudy (kendricklamar2014):

thanks

OpenStudy (helpblahblahblah):

anytime :)